a piece of wire that has been bent in the form of a semicircle including the bounding diameter is straightened and then bent in the form of a square.diameter of semicircle is 14cm.which has larger are semicircle or square?and by what difference.
ans.i have tried this way..area of circle(1/2 pi r square with 7 as radius.)
and taken the side of the square as 14 cm is it correct.what no should i use for side of the square?

Answer by mananth(16946)

diameter of semi circle = 14. so radius =7
length of the wire = 1/2* 2*pi*r +14
=22/7 *7
= 22+14 is the length of wire.
side of square = 36/4
=9
Area of semi circle 1/2 pi*r^2
1/2 *( 22/7 )* 7*7
= 77 sq. cm

Area of square = 9*9 = 81 sq.cm